A mixture contains 1 mole of helium (Cp = 2.5 R, Cv = 1.5 R) and 1 mole of hydrogen (Cp = 3.5 R, Cv = 2.5 R). Calculate the values of Cp, Cv and γ for the mixture.

Grade:10

1 Answers

Navjyot Kalra
9 years ago
Sol. Cp’ = 2.5 RCp” = 3.5 R Cv’ = 1.5 R Cv” = 2.5 R n base 1 = n base 2 = 1 mol (n base 1 + n base 2)CVdT = n base 1 C base vdT + n base 2 C base vdT ⇒ Cv = n base 1Cv’+ n base 2 Cv”/n base 1 + n base 2 = 1.5R + 2.5R/2 2R C base P = C base V + R = 2R + R = 3R γ = C base p/C base V = 3R/2R = 1.55